All systems are go

Don't agree with this? Please let us know

  • Live - CDN

  • HTTP - CDN

  • Control Panel

  • API
  • Cloud Storage

  • File Transcoding

  • Live Transcoding

  • Traffic Director

  • Managed DNS

  • Live Encoding as a Service

  • IPTV as a Service

Previous Incidents

[Resolved] Issue with upstream provider

This incident lasted 15 minutes.
  • Live - CDN
  • HTTP - CDN
Sat, 23 Sep 2017
09:51:05 UTC

We are currently experiencing issue with our upstream provider.For the time being, we will not be able to provide a stable and steady service. Will keep updating you about the status of the issue and we are always grateful for your kind support and cooperation.

10:01:02 UTC

The issue has been identified and a fix is being implemented.

10:06:05 UTC

This incident has been resolved.

[Complete] Scheduled APP server maintenance

This was a planned incident, it took 1 hour and 1 second.
  • Control Panel
  • API
Fri, 22 Sep 2017
22:08:10 UTC
Overview

We will be performing a system maintenance scheduled in our APP server .This maintenance will start at 03:30 UTC and will last a few minutes, during which none of our customers will be able to access their control panel and the API will be unavailable.There will not be any service impact on CDN level.

If you have any questions, please feel free to contact our Support Team: https://www.5centscdn.net/contact-us.html

  • Scheduled for Sat, 23 Sep 2017 3:30AM UTC
  • Expected to take 1 hour
Sat, 23 Sep 2017
03:30:02 UTC

The scheduled maintenance is now underway. We'll keep you updated on our progress.

04:30:02 UTC

The maintenance is now complete. Thanks for your patience.

[Resolved] Degraded Performance

This incident lasted 7 hours and 32 minutes.
  • Live - CDN
  • HTTP - CDN
Tue, 19 Sep 2017
16:55:18 UTC

We are currently experiencing degraded performance issue in 3 regions. We will update our status page as soon as more information becomes available.

This incident affects: Asia (NRT - Tokyo, Japan) and Europe (FRA - Frankfurt, Germany, LHR - London, United Kingdom).

17:45:32 UTC

We have identified the issue, currently 3 PoPs are affected. Our engineers are working around the clock, trying to resolve the issue as soon as possible.

18:03:52 UTC

A fix has been implemented and we are monitoring the results.

Wed, 20 Sep 2017
00:28:08 UTC

All services have been returned to a normal operating state.

No further notices from the past 7 days.